# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ 3.1.3 Robust Model Predictive Control
RMPC is designed to generate control actions that account for the possible realisations of uncertainties and disturbances over the prediction horizon. The previously discussed types of MPC optimise control actions based on a single set of plant predictions over the prediction horizon. In contrast, RMPC aims to optimise for a distribution of possible predictions, or for multiple samples drawn from said distribution. The distribution or samples correspond to all or particular ways in which the system may behave under different realisations of uncertainties and disturbances. Considering more than one prediction greatly increases the dimensionality and computational complexity of the underlying optimisation problem. However, there are formulations of RMPC that can do so efficiently by considering the statistical properties of uncertainties and disturbances, or by considering only a few relevant realisations of such values over a short horizon termed the robust horizon ($N_{r}$). Examples of the former formulations include min-max RMPC Witsenhausen (1968); Scokaert and Mayne (1998); Diehl and Bjornberg (2004) and Tube RMPC Mayne et al. (2005, 2011) while the latter formulation refers to multi stage RMPC (msRMPC) Lucia et al. (2013); Lucia (2014), also known as scenario RMPC. The plant models used in these types of MPC are the same as for LMPC and RMPC but contain additional terms and parameters that account for uncertainties and disturbances as in (3.1). Cost functions, on the other hand, are highly formulation-specific. As for NMPC, the computational cost of RMPC is a great motivation for the use of NNBO instead of conventional solvers.
@@ 3.2.1 Approximation via imitation learning
    Pure regression: When an NN is used to fit training data generated by sampling the original controller many times and the result is applied as is to replace the controller Kumar et al. (2018); Parisini and Zoppoli (1995); Lucia et al. (2020); Löhr et al. (2019); Lucia and Karg (2018); Bonzanini et al. (2021).
@@
    Steady state error compensation: When additional steps are taken to ensure that an approximation trained via pure regression does not suffer from steady state error or cause divergence from a steady state Åkesson et al. (2005); Chan et al. (2021); Li et al. (2022).
@@
    With feature engineering: When the training data generated by running the original controller is used to craft artificial features that serve as inputs to the NNBO approximation. The training still occurs via pure regression and the crafting of features can include the use of inputs that are not ever presented to the original MPC controller Drgoňa et al. (2018); Ortega and Camacho (1996); Vaupel et al. (2019); Nurbayeva et al. (2022); Karg and Lucia (2018); Lovelett et al. (2020); Parisini et al. (1998).
@@
    MPC inspired loss: When the regression error metric of the fitting problem is set to a function similar to the cost function of the original MPC controller Åkesson and Toivonen (2006); Drgoňa et al. (2022); Ahmed and Al-Dajani (1998).
@@
    Custom data generation: When the training data for pure regression is generated through more intricate manners than iterative evaluation of the original controller at random points. For instance, via evaluation of the original controller over trajectories in closed-loop form, or by accounting for state estimation errors, etc Cavagnari et al. (1999); Karg and Lucia (2019).
@@
    Primal-dual regressions: When the dual of the optimisation problem in the original controller is used or considered during NNBO training in an effort to exploit the properties of duality Wang et al. (2022).
@@ 3.2.2 Approximation with validation steps
These are approaches where trained approximations are put through tests to verify if specific performance requirements are met. Meeting said requirements must grant the approximation some form of guarantee, therefore, the standard practice of verifying the performance of a trained network on a validation or test set is excluded. From the works reviewed here, we have identified three sub categories that are relevant. Namely the cases where performance requirements are derived from results in statistics Zhang et al. (2020), probability Hertneck et al. (2018); Nubert et al. (2020); Karg et al. (2021), or ties to control theory Karg and Lucia (2020a, b); Fabiani and Goulart (2022); Pin et al. (2010, 2013).
@@ 3.2.3 Approximation with post-processing
These are frameworks which coerce the outputs of approximations generated via imitation learning in an effort to preserve some guarantee or improve performance. The coercing action can take several forms such as using the output of the approximation to hot start subsequent optimisation Chen et al. (2022); Löhr et al. (2020); Vaupel et al. (2020); Daosud et al. (2019), projecting the outputs onto sets that guarantee certain properties Chen et al. (2018); Bonzanini et al. (2020); Paulson and Mesbah (2020), or using the approximation itself to kickstart the training of an RL policy Karg and Lucia (2021). We do not include the simple case of output saturation here because it is frequent practice in all other categories.
@@ 4 Neural Network Based Optimisation for Linear Model Predictive Control
Åkesson et al. (2005) 1b
Train shallow NN to imitate LMPC using RMSE loss function, remove setpoint tracking error manually by zeroing NN outputs for setpoints.
None
2 1 1 20 20 4,000 - 5,000x Sim
Drgoňa et al. (2018) 1c
Train DNN with MSE loss to imitate MPC. DNN inputs defined via PCA and manual selection.
None
286 6 6 22 22 7.6x Sim
Kumar et al. (2018) 1a
Approximation of MPC with NN that combines LSTM and DNN. LSTM learns dependence on past actions while DNN learns dependence on current states.
None
4 1 1 Sim
Zhang et al. (2020) 2a
Fit DNN to primal and dual problems thereby enabling online constraint satisfaction check. If violations occur, call backup controller.
Probabilistic optimality with ability to check constraint satisfaction online.
4 3 3 3 3 10x Robot
Chen et al. (2022) 3a
Train DNN to initialise active set solver and thereby reduce iterations to convergence.
Recursive feasibility and asymptotic stability.
36 36 9 50 50 2x Sim
Chen et al. (2018) 3b
Project the outputs of an approximate DNN solver onto feasible output sets such that recursive feasibility is maintained. Train the DNN with policy gradient RL instead of supervised learning.
Recursive feasibility of inputs and states.
4 4 2 10 10 Sim
Karg and Lucia (2020a) 2c
Paper provides requirements for exact approximation of Explicit LMPC with DNN.
All EMPC guarantees for exact fits. Only constraint satisfaction for approximations.
4 1 1 10 10 Sim
Karg and Lucia (2020b) 2c
Verify input and state constraint satisfaction by solving a MILP that performs output range analysis on DNN ReLU approximation of solver, modify the approximation to behave like an LQR controller near the equilibrium point.
Input and state constraint satisfaction and stability.
2 1 1 3 3 Sim
@@
Fabiani and Goulart (2022) 2c
Derives conditions based on worst-case error and Lipschitz constant under which DNN ReLU approximations are stable. Formulates MILPs to calculate such quantities.
Optimality, stability, and recursive feasibility.
8 8 3 5 5 Sim
@@ 4.1 Methods without guarantees
The simplest way of generating an NNBO approximation of MPC is to train a NN to mimic the outputs of a conventional optimiser via imitation learning. This is a type of regression problem where the inputs of the network include some or all of the inputs of the optimiser, such as current system states and reference trajectories. The regressed outputs are the optimal control actions in the first step of the control horizon calculated by the conventional optimiser. This idea was first proposed by Parisini and Zoppoli Parisini and Zoppoli (1995) for an unconstrained receding horizon regulator with a nonlinear plant model. Later on Åkesson et al. Åkesson et al. (2005) applied this methodology to LMPC. They used a shallow NN with hyperbolic tangent activation and trained the network using the Levenberg-Marquardt (LM) algorithm on data generated by solving the LMPC on closed loop trajectories corresponding to step changes in setpoints. One of the issues of this approach is that due to approximation error, the outputs of the resulting NN can be non-zero at setpoints and when this occurs, the system starts to diverge from the setpoint. Åkesson et al. Åkesson et al. (2005) addressed this issue by manually zeroing the network outputs when all state derivatives are zero and the current output is equal to the desired setpoint. The effectiveness of the approach was demonstrated via simulation and the resulting approximate LMPC was 4,000 – 5,000 times faster than the original LMPC controller. Drgoňa et al. Drgoňa et al. (2018) showed that this approach could be extended to large problems with hundreds of state variables. In their work a Deep NN (DNN) was used instead of a shallow network and the inputs to the network were selected using expert knowledge of the system, Principal Component Analysis (PCA) and feature crafting. The study showed that the input features of the network can extend beyond what is available to the conventional optimiser in order to ease or increase the accuracy of the DNN regression. Similarly, Kumar et al. Kumar et al. (2018) showed that more advanced network architectures could be used to achieve the same purpose. Their framework combines a feedforward DNN with a network with Long Short-Term Memory (LSTM) units. The feedforward network learns the dependence of outputs on current states and the LSTM network learns the dependence of outputs on previous outputs similar to how a conventional solver is started from previous solutions. The networks are trained together and their output is fused with a weighted average. The authors showed through case studies that the proposed approach was a better regression than either network would produce independently. However, only a very small problem was considered, as shown in Table 1, and no details of computation time improvements with respect to the original controller were provided. Therefore, it is not clear if this method generalises well to larger problems or if it enables real time MPC, but the idea of using more complex neural units and network architectures is noteworthy.
@@
The common problem of approaches without guarantees is that safety of the controller cannot be ensured by design. Despite their differences, all aforementioned approaches boil down to solving a regression problem and although NNs are universal function approximators Barron (1993), exact regressions are impractical or intractable in practice and therefore regression error is unavoidable. With a sufficiently complex network and enough training data, it is possible to minimize the regression error substantially. However, as highlighted by Åkesson et al. Åkesson et al. (2005), this error can grow over time when the network is engaged in closed-loop control. Ultimately, the performance of the approximate LMPC controller may deviate significantly from that of the exact controller. Hence, approaches without guarantees require the use of add-on safety mechanism that run in parallel to the controller and perform functions such as control action saturation, constraint violation detection, instability detection, and emergency stopping. This arrangement is acceptable for some practical applications such as controlling inexpensive unmanned systems operating in low risk environments. However, for more critical systems the theoretical guarantees of MPC are desirable.
@@ 4.2 Methods with probabilistic guarantees
When a conventional optimisation solver converges to a solution, that solution is guaranteed to comply with all the constraints of the problem and be locally or globally optimal, within the tolerance limits stipulated by the settings of the solver. Zhang et al. Zhang et al. (2019, 2020) proposed a means to replicate this behaviour with high probability by training DNNs to imitate the output of conventional solvers for both the primal optimal control problem and its corresponding dual. Based on the properties of duality, the resulting primal and dual approximations can be used to conduct suboptimality and constraint satisfaction checks at runtime. If requirements are not met, a backup controller can be used instead of the calculated control actions. The study showed how the training problems can be formulated as a chance-constrained optimisation problem which can be approximated through randomisation. Thus, if the number of training samples is chosen adequately, both the primal and dual approximations can be trained to meet constraints and prescribed suboptimality thresholds with high probability. Statistical guarantees on the approximate randomised problem are purely dependent on the number of random samples used to solve the problem and the number of samples can be computed from scenario optimisation theory for convex problems, and from statistical learning theory for non-convex problems. The process of training a DNN approximation falls under the latter category (i.e. non-convex approximate chance-constrained problems). Since the calculated number of samples required to provide guarantees is generally very large, a heuristic can be adopted where a subset of samples is used for training and then the trained DNN is tested on all remaining samples. By following this framework one can train fast DNN based approximations of LMPC which comply with the input, output, and state constraints of the exact optimal control problem with high probability, and which yield results that meet user prescribed suboptimality bounds with high probability. Most importantly, the approach also enables the detection of instances where probabilistic guarantees have failed during operation. This can be used to trigger the activation of a backup control strategy, thereby ensuring the continuity of safe operation. The framework was demonstrated on an LVMPC integrated chassis control problem and the study reported a 10x speedup on the target hardware used in the automotive application. These results are very positive but it should be noted that the length of the horizon considered in the study was only three time steps and the underlying plant only had four states and three outputs. Yet, despite the small dimensions of the problem, the primal and dual DNN approximations had output dimensions 9 and 36, respectively. For larger plants and controllers with longer horizons, the ratio of inputs to outputs of the DNNs would be very low making the regression problem increasingly difficult and likely infeasible. Therefore, scalability of this method to larger problems is unlikely.
@@ 4.3 Methods with strict guarantees
Chen et al. Chen et al. (2022) proposed an explicit-implicit MPC approach where a DNN optimiser is used to warm start an active set solver. With this setup all MPC guarantees are preserved because a conventional solver is still used at every control iteration. The DNN optimiser is based on a primal formulation where the output of the network contains the primal variables for the entire prediction horizon and the loss function is based on the error between the values of the Lagrangian for the predicted and real solutions. By themselves the outputs of the network have no guarantees but because all primal variables are predicted, a simple algebraic check can be performed to assess the feasibility and suboptimality of the solution. If the solution is infeasible or suboptimal beyond some acceptable threshold then the output of the DNN is used to warm start an active set solver that coerces the solution to feasible and optimal values. Through simulation studies involving large plants and long horizons, Chen et al. Chen et al. (2022) showed that initialising the active set method with the solution found by the DNN significantly sped up the computation compared to the case where the active set method was cold started. However, cold starts are normally only required on the first iteration of MPC, on subsequent iterations warm starts can be performed using the previous solution. Therefore, this approach only benefits controllers that can already run in real time. Later on, Chen et al. Chen et al. (2018) proposed a different approach in which the outputs of a DNN approximation are projected onto feasible sets using Dykstra’s projection. This coercion preserves the recursive feasibility guarantees of MPC but requires significant overhead to perform the projection, and computation of the maximal control invariant set which is only feasible for some problems. For this study the authors demonstrated the method via simulation for a small plant and did not report on computation time improvements.
@@
Regarding the development of theoretical ties between MPC and NN approximations thereof, Karg and Lucia presented seminal results for approximations of EMPC Karg and Lucia (2020a) and LMPC  Karg and Lucia (2020b). In  Karg and Lucia (2020a), the authors derived the requirements that a DNN with Rectified Linear Unit (ReLU) activation functions needs to satisfy to exactly fit an EMPC policy. They also suggested methods for quantifying the approximation error of networks that did not meet these requirements. Then, in Karg and Lucia (2020b), the authors started developing similar theory for LMPC. Specifically, they designed a MILP that can be applied to each output of a DNN ReLU approximation of LMPC to perform output range analysis and determine if the output in question can violate its constraints during closed loop operation. In practice generating an approximation that passes the MILP checks for all outputs can be difficult and therefore the authors also proposed a method to modify trained approximations via optimisation and make them behave like an LQR controller near the equilibrium region. In both studies Karg and Lucia (2020a, b), the authors proved the effectiveness of their methods via simulation with small plants with a single input and a controller with a short horizon as shown in Table 1. For LMPC controllers of that size, conventional optimisers can run in real time at thousands of $Hz$ and because the methods proposed by Karg and Lucia have not been demonstrated on larger problems their scalability is unclear. Most recently, Fabiani and Goulart Fabiani and Goulart (2022) built on the foundations by Karg and Lucia Karg and Lucia (2020a) and proposed an algorithm involving two MILPs to fit a minimal complexity DNN ReLU that approximates a given LMPC controller and preserves strict stability and constraint violation guarantees. The output of the first MILP is the upper bound on approximation error of the DNN in terms of the difference in computed optimal inputs with respect to a real optimiser. The solution of the second MILP is the Lipschitz constant associated to the approximation error. If the constant is below a certain threshold then the DNN outputs are guaranteed to ensure closed-loop stability around the origin of the plant. Importantly, this framework always requires solving two MILPs while the method by Karg and Lucia Karg and Lucia (2020b) requires solving a MILP per network output. Fabiani and Goulart Fabiani and Goulart (2022) also noted that while MI optimisation typically has poor scalability with problem size, their numerical experiments showed weak dependence of MILP solution times on number of plant states and strong dependence on number of inputs. However, their experiments used a prediction horizon of only five steps and the scalability of the method with prediction horizon length was not discussed. Overall, good progress has been made in developing NN based approximations of LMPC with strict guarantees but such methods have only been shown to work in small problems and their scalability has not been explored in depth.

### v3

@@ 3.1.3 Robust Model Predictive Control
RMPC is designed to generate control actions that account for the possible realisations of uncertainties and disturbances over the prediction horizon. The previously discussed types of MPC optimise control actions based on a single set of plant predictions over the prediction horizon. In contrast, RMPC aims to optimise for a distribution of possible predictions, or for multiple samples drawn from said distribution. The distribution or samples correspond to all or particular ways in which the system may behave under different realisations of uncertainties and disturbances. Considering more than one prediction greatly increases the dimensionality and computational complexity of the underlying optimisation problem. However, there are formulations of RMPC that can do so efficiently by considering the statistical properties of uncertainties and disturbances, or by considering only a few relevant realisations of such values over a short horizon termed the robust horizon ($N_{r}$). Examples of the former formulations include min-max RMPC (Witsenhausen, 1968; Scokaert and Mayne, 1998; Diehl and Bjornberg, 2004) and Tube RMPC (Mayne et al., 2005, 2011) while the latter formulation refers to multi stage RMPC (msRMPC) (Lucia et al., 2013; Lucia, 2014), also known as scenario RMPC. The plant models used in these types of MPC are the same as for LMPC and RMPC but contain additional terms and parameters that account for uncertainties and disturbances as in (1). Cost functions, on the other hand, are highly formulation-specific. As for NMPC, the computational cost of RMPC is a great motivation for the use of NNBO instead of conventional solvers.
@@ 3.2.1 Approximation via imitation learning
    Pure regression: When an NN is used to fit training data generated by sampling the original controller many times and the result is applied as is to replace the controller (Kumar et al., 2018; Parisini and Zoppoli, 1995; Lucia et al., 2020; Löhr et al., 2019; Lucia and Karg, 2018; Bonzanini et al., 2021).
@@
    Steady state error compensation: When additional steps are taken to ensure that an approximation trained via pure regression does not suffer from steady state error or cause divergence from a steady state (Åkesson et al., 2005; Chan et al., 2021; Li et al., 2022).
@@
    With feature engineering: When the training data generated by running the original controller is used to craft artificial features that serve as inputs to the NNBO approximation. The training still occurs via pure regression and the crafting of features can include the use of inputs that are not ever presented to the original MPC controller (Drgoňa et al., 2018; Ortega and Camacho, 1996; Vaupel et al., 2019; Nurbayeva et al., 2022; Karg and Lucia, 2018; Lovelett et al., 2020; Parisini et al., 1998).
@@
    MPC inspired loss: When the regression error metric of the fitting problem is set to a function similar to the cost function of the original MPC controller (Åkesson and Toivonen, 2006; Drgoňa et al., 2022; Ahmed and Al-Dajani, 1998).
@@
    Custom data generation: When the training data for pure regression is generated through more intricate manners than iterative evaluation of the original controller at random points. For instance, via evaluation of the original controller over trajectories in closed-loop form, or by accounting for state estimation errors, etc (Cavagnari et al., 1999; Karg and Lucia, 2019).
@@
    Primal-dual regressions: When the dual of the optimisation problem in the original controller is used or considered during NNBO training in an effort to exploit the properties of duality (Wang et al., 2022).
@@ 3.2.2 Approximation with validation steps
These are approaches where trained approximations are put through tests to verify if specific performance requirements are met. Meeting said requirements must grant the approximation some form of guarantee, therefore, the standard practice of verifying the performance of a trained network on a validation or test set is excluded. From the works reviewed here, we have identified three sub categories that are relevant. Namely the cases where performance requirements are derived from results in statistics (Zhang et al., 2020), probability (Hertneck et al., 2018; Nubert et al., 2020; Karg et al., 2021), or ties to control theory (Karg and Lucia, 2020a, b; Fabiani and Goulart, 2022; Pin et al., 2010, 2013).
@@ 3.2.3 Approximation with post-processing
These are frameworks which coerce the outputs of approximations generated via imitation learning in an effort to preserve some guarantee or improve performance. The coercing action can take several forms such as using the output of the approximation to hot start subsequent optimisation (Chen et al., 2022; Löhr et al., 2020; Vaupel et al., 2020; Daosud et al., 2019), projecting the outputs onto sets that guarantee certain properties (Chen et al., 2018; Bonzanini et al., 2020; Paulson and Mesbah, 2020), or using the approximation itself to kickstart the training of an RL policy (Karg and Lucia, 2021). We do not include the simple case of output saturation here because it is frequent practice in all other categories.
@@ 4 Neural Network Based Optimisation for Linear Model Predictive Control
Åkesson et al. (2005) 1b Train shallow NN to imitate LMPC using RMSE loss function, remove setpoint tracking error manually by zeroing NN outputs for setpoints. None 2 1 1 20 20 4,000 - 5,000x Sim
Drgoňa et al. (2018) 1c Train DNN with MSE loss to imitate MPC. DNN inputs defined via PCA and manual selection. None 286 6 6 22 22 7.6x Sim
Kumar et al. (2018) 1a Approximation of MPC with NN that combines LSTM and DNN. LSTM learns dependence on past actions while DNN learns dependence on current states. None 4 1 1 Sim
Zhang et al. (2020) 2a Fit DNN to primal and dual problems thereby enabling online constraint satisfaction check. If violations occur, call backup controller. Probabilistic optimality with ability to check constraint satisfaction online. 4 3 3 3 3 10x Robot
Chen et al. (2022) 3a Train DNN to initialise active set solver and thereby reduce iterations to convergence. Recursive feasibility and asymptotic stability. 36 36 9 50 50 2x Sim
Chen et al. (2018) 3b Project the outputs of an approximate DNN solver onto feasible output sets such that recursive feasibility is maintained. Train the DNN with policy gradient RL instead of supervised learning. Recursive feasibility of inputs and states. 4 4 2 10 10 Sim
Karg and Lucia (2020a) 2c Paper provides requirements for exact approximation of Explicit LMPC with DNN. All EMPC guarantees for exact fits. Only constraint satisfaction for approximations. 4 1 1 10 10 Sim
Karg and Lucia (2020b) 2c Verify input and state constraint satisfaction by solving a MILP that performs output range analysis on DNN ReLU approximation of solver, modify the approximation to behave like an LQR controller near the equilibrium point. Input and state constraint satisfaction and stability. 2 1 1 3 3 Sim
@@
Fabiani and Goulart (2022) 2c Derives conditions based on worst-case error and Lipschitz constant under which DNN ReLU approximations are stable. Formulates MILPs to calculate such quantities. Optimality, stability, and recursive feasibility. 8 8 3 5 5 Sim
@@ 4.1 Methods without guarantees
The simplest way of generating an NNBO approximation of MPC is to train a NN to mimic the outputs of a conventional optimiser via imitation learning. This is a type of regression problem where the inputs of the network include some or all of the inputs of the optimiser, such as current system states and reference trajectories. The regressed outputs are the optimal control actions in the first step of the control horizon calculated by the conventional optimiser. This idea was first proposed by Parisini and Zoppoli (1995) for an unconstrained receding horizon regulator with a nonlinear plant model. Later on, Åkesson et al. (2005) applied this methodology to LMPC. They used a shallow NN with hyperbolic tangent activation and trained the network using the Levenberg-Marquardt (LM) algorithm on data generated by solving the LMPC on closed loop trajectories corresponding to step changes in setpoints. One of the issues of this approach is that due to approximation error, the outputs of the resulting NN can be non-zero at setpoints and when this occurs, the system starts to diverge from the setpoint. Åkesson et al. (2005) addressed this issue by manually zeroing the network outputs when all state derivatives are zero and the current output is equal to the desired setpoint. The effectiveness of the approach was demonstrated via simulation and the resulting approximate LMPC was 4,000 – 5,000 times faster than the original LMPC controller. Drgoňa et al. (2018) showed that this approach could be extended to large problems with hundreds of state variables. In their work a Deep NN (DNN) was used instead of a shallow network and the inputs to the network were selected using expert knowledge of the system, Principal Component Analysis (PCA) and feature crafting. The study showed that the input features of the network can extend beyond what is available to the conventional optimiser in order to ease or increase the accuracy of the DNN regression. Similarly, Kumar et al. (2018) showed that more advanced network architectures could be used to achieve the same purpose. Their framework combines a feedforward DNN with a network with Long Short-Term Memory (LSTM) units. The feedforward network learns the dependence of outputs on current states and the LSTM network learns the dependence of outputs on previous outputs similar to how a conventional solver is started from previous solutions. The networks are trained together and their output is fused with a weighted average. The authors showed through case studies that the proposed approach was a better regression than either network would produce independently. However, only a very small problem was considered, as shown in Table 1, and no details of computation time improvements with respect to the original controller were provided. Therefore, it is not clear if this method generalises well to larger problems or if it enables real time MPC, but the idea of using more complex neural units and network architectures is noteworthy.
@@
The common problem of approaches without guarantees is that safety of the controller cannot be ensured by design. Despite their differences, all aforementioned approaches boil down to solving a regression problem and although NNs are universal function approximators (Barron, 1993), exact regressions are impractical or intractable in practice and therefore regression error is unavoidable. With a sufficiently complex network and enough training data, it is possible to minimize the regression error substantially. However, as highlighted by Åkesson et al. (2005), this error can grow over time when the network is engaged in closed-loop control. Ultimately, the performance of the approximate LMPC controller may deviate significantly from that of the exact controller. Hence, approaches without guarantees require the use of add-on safety mechanism that run in parallel to the controller and perform functions such as control action saturation, constraint violation detection, instability detection, and emergency stopping. This arrangement is acceptable for some practical applications such as controlling inexpensive unmanned systems operating in low risk environments. However, for more critical systems the theoretical guarantees of MPC are desirable.
@@ 4.2 Methods with probabilistic guarantees
When a conventional optimisation solver converges to a solution, that solution is guaranteed to comply with all the constraints of the problem and be locally or globally optimal, within the tolerance limits stipulated by the settings of the solver. Zhang et al. (2019, 2020) proposed a means to replicate this behaviour with high probability by training DNNs to imitate the output of conventional solvers for both the primal optimal control problem and its corresponding dual. Based on the properties of duality, the resulting primal and dual approximations can be used to conduct suboptimality and constraint satisfaction checks at runtime. If requirements are not met, a backup controller can be used instead of the calculated control actions. The study showed how the training problems can be formulated as a chance-constrained optimisation problem which can be approximated through randomisation. Thus, if the number of training samples is chosen adequately, both the primal and dual approximations can be trained to meet constraints and prescribed suboptimality thresholds with high probability. Statistical guarantees on the approximate randomised problem are purely dependent on the number of random samples used to solve the problem and the number of samples can be computed from scenario optimisation theory for convex problems, and from statistical learning theory for non-convex problems. The process of training a DNN approximation falls under the latter category (i.e. non-convex approximate chance-constrained problems). Since the calculated number of samples required to provide guarantees is generally very large, a heuristic can be adopted where a subset of samples is used for training and then the trained DNN is tested on all remaining samples. By following this framework one can train fast DNN based approximations of LMPC which comply with the input, output, and state constraints of the exact optimal control problem with high probability, and which yield results that meet user prescribed suboptimality bounds with high probability. Most importantly, the approach also enables the detection of instances where probabilistic guarantees have failed during operation. This can be used to trigger the activation of a backup control strategy, thereby ensuring the continuity of safe operation. The framework was demonstrated on an LVMPC integrated chassis control problem and the study reported a 10x speedup on the target hardware used in the automotive application. These results are very positive but it should be noted that the length of the horizon considered in the study was only three time steps and the underlying plant only had four states and three outputs. Yet, despite the small dimensions of the problem, the primal and dual DNN approximations had output dimensions 9 and 36, respectively. For larger plants and controllers with longer horizons, the ratio of inputs to outputs of the DNNs would be very low making the regression problem increasingly difficult and likely infeasible. Therefore, scalability of this method to larger problems is unlikely.
@@ 4.3 Methods with strict guarantees
Chen et al. (2022) proposed an explicit-implicit MPC approach where a DNN optimiser is used to warm start an active set solver. With this setup all MPC guarantees are preserved because a conventional solver is still used at every control iteration. The DNN optimiser is based on a primal formulation where the output of the network contains the primal variables for the entire prediction horizon and the loss function is based on the error between the values of the Lagrangian for the predicted and real solutions. By themselves the outputs of the network have no guarantees but because all primal variables are predicted, a simple algebraic check can be performed to assess the feasibility and suboptimality of the solution. If the solution is infeasible or suboptimal beyond some acceptable threshold then the output of the DNN is used to warm start an active set solver that coerces the solution to feasible and optimal values. Through simulation studies involving large plants and long horizons, Chen et al. (2022) showed that initialising the active set method with the solution found by the DNN significantly sped up the computation compared to the case where the active set method was cold started. However, cold starts are normally only required on the first iteration of MPC, on subsequent iterations warm starts can be performed using the previous solution. Therefore, this approach only benefits controllers that can already run in real time. Later on, Chen et al. (2018) proposed a different approach in which the outputs of a DNN approximation are projected onto feasible sets using Dykstra’s projection. This coercion preserves the recursive feasibility guarantees of MPC but requires significant overhead to perform the projection, and computation of the maximal control invariant set which is only feasible for some problems. For this study the authors demonstrated the method via simulation for a small plant and did not report on computation time improvements.
@@
Regarding the development of theoretical ties between MPC and NN approximations thereof, Karg and Lucia presented seminal results for approximations of EMPC (Karg and Lucia, 2020a) and LMPC (Karg and Lucia, 2020b). In Karg and Lucia (2020a), the authors derived the requirements that a DNN with Rectified Linear Unit (ReLU) activation functions needs to satisfy to exactly fit an EMPC policy. They also suggested methods for quantifying the approximation error of networks that did not meet these requirements. Then, in Karg and Lucia (2020b), the authors started developing similar theory for LMPC. Specifically, they designed a MILP that can be applied to each output of a DNN ReLU approximation of LMPC to perform output range analysis and determine if the output in question can violate its constraints during closed loop operation. In practice generating an approximation that passes the MILP checks for all outputs can be difficult and therefore the authors also proposed a method to modify trained approximations via optimisation and make them behave like an LQR controller near the equilibrium region. In both studies, Karg and Lucia (2020a, b) proved the effectiveness of their methods via simulation with small plants with a single input and a controller with a short horizon as shown in Table 1. For LMPC controllers of that size, conventional optimisers can run in real time at thousands of $Hz$ and because the methods proposed by Karg and Lucia have not been demonstrated on larger problems their scalability is unclear. Most recently, Fabiani and Goulart (2022) built on the foundations by Karg and Lucia (2020a) and proposed an algorithm involving two MILPs to fit a minimal complexity DNN ReLU that approximates a given LMPC controller and preserves strict stability and constraint violation guarantees. The output of the first MILP is the upper bound on approximation error of the DNN in terms of the difference in computed optimal inputs with respect to a real optimiser. The solution of the second MILP is the Lipschitz constant associated to the approximation error. If the constant is below a certain threshold then the DNN outputs are guaranteed to ensure closed-loop stability around the origin of the plant. Importantly, this framework always requires solving two MILPs while the method by Karg and Lucia (2020b) requires solving a MILP per network output. Fabiani and Goulart (2022) also noted that while MI optimisation typically has poor scalability with problem size, their numerical experiments showed weak dependence of MILP solution times on number of plant states and strong dependence on number of inputs. However, their experiments used a prediction horizon of only five steps and the scalability of the method with prediction horizon length was not discussed. Overall, good progress has been made in developing NN based approximations of LMPC with strict guarantees but such methods have only been shown to work in small problems and their scalability has not been explored in depth.